OpenStudy (anonymous):

let X and Y be two continuous random variables with joint density,can someone help

OpenStudy (anonymous):

f(x,y)=\[\left\{ 4xy,0<x<1,0<y<1 \right\} \] find the joint density of w=x^2 and z =xy

OpenStudy (anonymous):

can u help mathdood

OpenStudy (anonymous):

Haha no I am just here to learn. If by some miracle I figure it out before someone else gets here then I would be happy to help lol.

OpenStudy (anonymous):

can u plz help zarkon

OpenStudy (zarkon):

this is just a change of variables....review that...look up Jacobian